# Beyond Fast and Slow: Cognitive-Inspired El-astic Reasoning for Large Language Models

### v1

@@ 1 Introduction
Large language models (LLMs), such as ChatGPT (openai2023gpt) and DeepSeek (guo2025deepseek), have achieved impressive results on many tasks, including multi-turn dialogue (stark2023dobby) and embodied intelligence (mu2023embodiedgpt). However, as model size and the number of inference tokens increase, the computational resources required for inference grow substantially, creating a major bottleneck for real-world applications. Meanwhile, user queries vary widely in complexity, from straightforward fact-based questions to multi-hop reasoning tasks, and in some cases, even require external tool invocation. This diversity makes traditional LLM reasoning approaches, rooted in the dual-process theory of fast (System 1) and slow (System 2) thinking, face critical limitations in handling all types of queries efficiently and effectively (li2025system). Consequently, it is crucial to dynamically allocate reasoning strategies based on query complexity in practical applications.
@@
Unfortunately, existing LLMs typically apply a uniform reasoning process regardless of query complexity (aggarwal2025l1). This one-size-fits-all reasoning strategy risks either wasting computation on trivial inputs or inadequately handling more demanding queries. Achieving flexible and efficient reasoning requires addressing two key challenges: 1) Unforeseen query difficulty: The true complexity of an incoming query is often not observable in advance, making it difficult to allocate computational resources dynamically and appropriately. 2) Cost–quality trade-off: Larger language models generally yield higher accuracy but incur substantially greater compute costs, forcing a careful balance between performance and efficiency along the Pareto frontier.
@@
Recently, several attempts (jiang2023llm; dong2024self; du2023improving; ong2025routellm; yang2025towards) have been proposed to tailor reasoning strategies to downstream task demands, which can be broadly divided into the following categories: 1) LLM ensemble methods (jiang2023llm; dong2024self; du2023improving) often combine outputs from multiple candidate models to boost accuracy. However, each input must typically be processed by all models in the ensemble, leading to substantial computational overhead. 2) Test-time scaling methods (muennighoff2025s1; yang2025towards; snell2024scaling; aggarwal2025l1)adapt reasoning costs based on the estimated difficulty of inputs, for instance by adjusting the length of chain-of-thought (CoT) reasoning or employing early-exit mechanisms. While more efficient, these methods often struggle to assess difficulty accurately for all queries and lack adaptive mechanisms for invoking external tools. As a result, they fall short in handling complex tasks requiring access to additional knowledge sources, limiting their flexibility and extensibility in real-world applications.
@@
To address these limitations, we propose the Cognitive-Inspired Elastic Reasoning (CogER) framework for efficient scaling of language model reasoning. This framework dynamically selects the most suitable processing mode for each query based on its complexity. Specifically, inspired by Bloom’s Taxonomy (bloom1956taxonomy), we first categorize incoming queries into four complexity levels ($L_{1}-L_{4}$), each associated with a tailored reasoning strategy, thereby mitigating the challenge of unforeseen query difficulty. Then, we model the strategy selection process as a Markov Decision Process (MDP), in which a CogER-Agent chooses one of four actions (No Think, Think, Extend, or Delegate) to process each query, based on the predicted complexity level. To guide the training of this agent, we design a reward function that explicitly balances computational cost against output quality, ensuring that each query receives only the computational resources commensurate with its complexity. Finally, for $L_{4}$ queries that require external knowledge, we introduce Cognitive Tool-Assisted Reasoning (CoTool), enabling the LLM to autonomously invoke external tools at appropriate points within its chain-of-thought, enabling flexible and knowledge-augmented reasoning.
@@ 2 Related Work
Large language models (LLMs) ensemble methods (chen2025harnessing) aim to combine multiple models to leverage their complementary strengths. Existing approaches can be categorized into three paradigms based on integration timing: ensemble-before-inference, ensemble-during-inference, and ensemble-after-inference. Ensemble‐before‐inference methods (lu2024routing; dinghybrid; srivatsa2024harnessing; lu2024blending) first apply a routing mechanism, either pretrained on custom data or trained on the fly, to dispatch each query to the most suitable, specialized model, thereby enabling more cost‐efficient inference. Ensemble-during-inference methods (huang2024ensemble; xu2025hit; park2024ensembling) combine outputs from multiple models at different levels of granularity, including the token level, span level, and reasoning-step level, and then merge the resulting text segments back into the decoding context to iteratively refine the output. Ensemble-after-inference methods (park2024ensembling; hu2025dynamic; du2023improving) generate complete responses independently from each candidate LLM and then consolidate them via ranking, majority voting, or fitness scoring to select the highest-quality output for final delivery. In contrast, our CogER learns a lightweight policy network to dispatch each query to a single, optimal inference action, $\mathrm{No\_Think}$, $\mathrm{Think}$, $\mathrm{Extend}$, $\mathrm{Delegate}$, within an MDP, thereby enabling fine-grained per-query adaptation.
Test-Time Scaling (TTS) methods optimize computational resource allocation during inference through adaptive reasoning depth control. muennighoff2025s1 propose budget forcing, a technique to regulate the computation of test time by prematurely stopping the CoT of the model or extending it by repeated insertion of the token ‘wait’ when the model attempts to terminate generation. aggarwal2025l1 propose LCPO, a straightforward reinforcement learning (RL) approach designed to maximize accuracy while respecting user‑specified length constraints. yang2025towards propose Thinking-Optimal Scaling, which trains the model on seed examples with varied response lengths to learn appropriate reasoning efforts and then self-improves by selecting the shortest correct responses on new tasks. In contrast to TTS methods that only adjust reasoning depth based on coarse difficulty estimates, CogER selects among diverse reasoning modes and seamlessly integrates external tool usage, achieving more versatile resource allocation.
@@ 3 Problem Statement and Motivation
Motivation. In real-world applications, user questions exhibit a wide range of complexity. For example, some queries can be answered in a single step, whereas others require deep, multi-step reasoning or integration of external information sources. However, existing LLMs apply the same reasoning procedure to every query with a high computational cost, which may lead to wasted resources on simple queries and poor performance on complex ones (sui2025stop; hu2025dynamic).
@@ 4.1 Query Complexity Classification
To efficiently allocate reasoning strategies based on the diverse computational requirements of different queries, we draw inspiration from Bloom’s Taxonomy (bloom1956taxonomy) to classify queries by their cognitive demand. Specifically, we define four levels of query complexity, denoted as $L_{1}$, $L_{2}$, $L_{3}$, and $L_{4}$, with each level representing an increasing degree of reasoning depth and computational demand, as illustrated in Figure 1(a). The details of each level are as follows:
@@ 4.2 Cognitive-Inspired Elastic Reasoning as Markov Decision Process
  • $L_{2}$-level: $\mathrm{Think}$. Apply a chain-of-thought strategy (wei2022chain) using a moderately sized LLM to produce a concise reasoning trail.
@@ 4.5 Training with Group Relative Policy Optimization
We adopt the Group Relative Policy Optimization (GRPO) (shao2024deepseekmath; guo2025deepseek) to optimize the parameters $\theta$ of the CogER-Agent due to its superior stability and sample-efficiency.
@@ 5 Experiments
Datasets and Metrics. To train the CogER-Agent, we construct the Reasoning‑Training dataset by randomly sampling 2,000 examples from each of four heterogeneous benchmarks: GSM8K (cobbe2021training), MATH (hendrycks2021measuring), CommonsenseQA (talmor2019commonsenseqa), and MedQA (jin2021disease). This unified training set exposes the agent to a wide spectrum of reasoning challenges, from arithmetic word problems to domain‑specific medical questions. For evaluation, we consider both In‑Domain (ID) and Out‑of‑Domain (OOD) settings. ID performance is evaluated on the official test splits of GSM8K, MATH-500, CommonsenseQA, and MedQA, whereas OOD generalization is measured on MAWPS (koncel2016mawps) and CollegeMath (tang2024mathscale), which are not included in the mixed training set used to fine-tune our CogER-Agent. More details in App. C. We report Exact Match ($EM$) as the metric across all datasets, and record the average parameters (Param.) and Latency used during testing to reflect computational cost.
Baselines. We employ LLMs with varying sizes and architectures, including Qwen2.5-Math-72B-Instruct (yang2024qwen2math), DeepSeek-R1 (guo2025deepseek), DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-14B, and DeepSeek-R1-Distill-Qwen-32B. We also compare against Test-Time Scaling (TTS) methods, including S1 (muennighoff2025s1), L1 (aggarwal2025l1), and ReasonFlux (yang2025reasonflux).
Implementation Details. In our CogER framework, Qwen2.5-7B-Instruct (yang2024qwen2) serves as the CogER-Agent, which assigns queries to appropriate reasoning modules based on their estimated complexity, and also handles all $L_{1}$‑level queries directly. Queries classified as $L_{2}$‑level are escalated to Qwen2.5‑32B‑Instruct for moderate multi‑step reasoning, while $L_{3}$‑level queries are processed by QwQ‑32B (qwq32b) to support deeper CoT generation. For the most demanding $L_{4}$‑level queries, we invoke our CoTool, whereby QwQ‑32B (qwq32b) autonomously issues external API calls to enrich its reasoning process. We uniformly capped the generation length at max_token = 8192 for all LLMs. Furthermore, all components are optimized using the AdamW optimizer with a batch size of $24\times 3$ and a learning rate of $5\times 10^{-5}$. The group size $G$ in Eqn. (10) is set to 12. The CogER-Agent is fine-tuned via LoRA with a rank of $r=16$, while all other hyperparameters follow the default settings from the Open-R1 configuration (openr1).
@@ A.1 Large Language Models
Recent advancements in Natural Language Processing (NLP) have culminated in the development of highly capable Large language models (LLMs) (hutest; wang2025enhancing). These models are predominantly characterized by their utilization of the Transformer architecture (vaswani2017attention) and are pre-trained on vast quantities of textual data. Progressive scaling of model parameters and training datasets has allowed these LLMs to exhibit emergent capabilities (wei2022emergent), demonstrating remarkable proficiency in the handling of a variety of complex tasks. Such tasks include, but are not limited to, high-fidelity question answering (shao2023prompting), code generation (chen2021evaluating), and intermediate-step reasoning (wei2022chain). Consequently, LLMs have exerted a profound influence on the Artificial Intelligence (AI) community, catalyzing a reevaluation of the prospects for Artificial General Intelligence (AGI) (zhao2023survey). Predicated on their foundational Transformer architecture, extant Large language models can be broadly classified into three primary categories:
@@
Encoder-only models. Encoder-only models, alternatively designated as auto-encoding architectures, are exclusively constructed from encoder modules originating from the Transformer architecture. In such configurations, input data is processed sequentially across multiple layers, facilitating the progressive extraction and encoding of information. A paradigmatic example of this category is BERT (Bidirectional Encoder Representations from Transformers) (devlin2018bert), developed by Google. BERT functions as a language representation model employing bidirectional Transformer encoders. It was pre-trained on a corpus consisting of the BooksCorpus (zhu2015aligning) (comprising approximately 800 million words) and English Wikipedia (approximately 2.5 billion words). This pre-training enabled BERT to attain a score of 80.5% on the General Language Understanding Evaluation (GLUE) benchmark and an accuracy of 86.7% on the Multi-Genre Natural Language Inference (MultiNLI) task. Many subsequent encoder-only models are predominantly variants of BERT, including RoBERTa by Meta (liu2019roberta) and DeBERTa by Microsoft (he2020deberta).
@@
Decoder-only models. This category of models is exclusively constructed using decoder modules from the Transformer architecture. Decoder-only models typically implement an auto-regressive mechanism, whereby output sequences are generated on a token-by-token basis. The generation of each token by the decoder is contingent upon the tokens previously generated. Seminal examples in this category include the Generative Pre-trained Transformer (GPT) series (openai2023gpt), developed by OpenAI. Illustratively, GPT-3 comprises numerous Transformer decoder layers, featuring up to 175 billion parameters, which established it as one of the most substantial language models at its introduction. This model was pre-trained on a corpus of approximately 300 billion tokens derived from sources such as Common Crawl (raffel2020exploring), WebText2, Books1, Books2, and Wikipedia. GPT-3 has demonstrated potent zero-shot and few-shot learning capabilities across a diverse array of language tasks. Beyond the GPT series, a multitude of other decoder-only models have emerged, including OPT (zhang2022opt), LLaMA (touvron2023llama), and Llama 2 (touvron2023llama2) from Meta; PaLM (chowdhery2023palm) and PaLM 2 (anil2023palm) from Google; and BLOOM (workshop2022bloom) from the BigScience initiative. Furthermore, models such as Qwen 2.5 by Alibaba (yang2024qwen2), alongside DeepSeek LLM (bi2024deepseek) and DeepSeek Coder (guo2024deepseek) by DeepSeek AI, continue to advance the state-of-the-art in language comprehension, multilingual processing, and domain-specific generation.
@@
Encoder-decoder models. This architectural class integrates both encoder and decoder modules from the Transformer framework. Such models aim to amalgamate the respective strengths of the aforementioned architectures, thereby effectively addressing tasks that necessitate comprehensive input understanding coupled with the generation of extended output sequences. Prominent extant encoder-decoder models encompass GLM from Tsinghua University (du2021glm); T5 (raffel2020exploring), FLAN-T5 (chung2024scaling), and UL2 (tay2022ul2) from Google; and BART (lewis2019bart) from Meta. For instance, the GLM model employs an autoregressive blank infilling objective. This methodology is designed to tackle three fundamental challenges in NLP: natural language understanding (NLU), unconditional text generation, and conditional text generation. With a reported maximum of 130 billion parameters, GLM was pre-trained on a corpus including BookCorpus (tay2022ul2) and Wikipedia. GLM has demonstrated superior performance over BERT on the SuperGLUE benchmark, exhibiting an improvement of 4.6%–5.0%. Moreover, it significantly surpasses FLAN-T5 in both NLU and generation tasks, even when utilizing fewer parameters and less training data.

### v2

@@ 1 Introduction
Large language models (LLMs), such as ChatGPT (OpenAI, 2023) and DeepSeek (DeepSeek-AI, 2025), have achieved impressive results on many tasks, including multi-turn dialogue (Stark et al., 2024) and embodied intelligence (Mu et al., 2023). However, as model size and the number of inference tokens increase, the computational resources required for inference grow substantially, creating a major bottleneck for real-world applications. Meanwhile, user queries vary widely in complexity, from straightforward fact-based questions to multi-hop reasoning tasks, and in some cases, even require external tool invocation. This diversity makes traditional LLM reasoning approaches, rooted in the dual-process theory of fast (System 1) and slow (System 2) thinking, face critical limitations in handling all types of queries efficiently and effectively (Li et al., 2025). Consequently, it is crucial to dynamically allocate reasoning strategies based on query complexity in practical applications.
@@
Unfortunately, existing LLMs typically apply a uniform reasoning process regardless of query complexity (Aggarwal and Welleck, 2025). This one-size-fits-all reasoning strategy risks either wasting computation on trivial inputs or inadequately handling more demanding queries. Achieving flexible and efficient reasoning requires addressing two key challenges: 1) Unforeseen query difficulty: The true complexity of an incoming query is often not observable in advance, making it difficult to allocate computational resources dynamically and appropriately. 2) Cost–quality trade-off: Larger language models generally yield higher accuracy but incur substantially greater compute costs, forcing a careful balance between performance and efficiency along the Pareto frontier.
@@
Recently, several attempts (Jiang et al., 2023; Dong et al., 2024; Du et al., 2024; Ong et al., 2025; Yang et al., 2025b) have been proposed to tailor reasoning strategies to downstream task demands, which can be broadly divided into the following categories: 1) LLM ensemble methods (Jiang et al., 2023; Dong et al., 2024; Du et al., 2024) often combine outputs from multiple candidate models to boost accuracy. However, each input must typically be processed by all models in the ensemble, leading to substantial computational overhead. 2) Test-time scaling methods (Muennighoff et al., 2025; Yang et al., 2025b; Snell et al., 2024; Aggarwal and Welleck, 2025)adapt reasoning costs based on the estimated difficulty of inputs, for instance by adjusting the length of chain-of-thought (CoT) reasoning or employing early-exit mechanisms. While more efficient, these methods often struggle to assess difficulty accurately for all queries and lack adaptive mechanisms for invoking external tools. As a result, they fall short in handling complex tasks requiring access to additional knowledge sources, limiting their flexibility and extensibility in real-world applications.
@@
To address these limitations, we propose the Cognitive-Inspired Elastic Reasoning (CogER) framework for efficient scaling of language model reasoning. This framework dynamically selects the most suitable processing mode for each query based on its complexity. Specifically, inspired by Bloom’s Taxonomy (Bloom et al., 1956), we first categorize incoming queries into four complexity levels ($L_{1}-L_{4}$), each associated with a tailored reasoning strategy, thereby mitigating the challenge of unforeseen query difficulty. Then, we model the strategy selection process as a Markov Decision Process (MDP), in which a CogER-Agent chooses one of four actions (No Think, Think, Extend, or Delegate) to process each query, based on the predicted complexity level. To guide the training of this agent, we design a reward function that explicitly balances computational cost against output quality, ensuring that each query receives only the computational resources commensurate with its complexity. Finally, for $L_{4}$ queries that require external knowledge, we introduce Cognitive Tool-Assisted Reasoning (CoTool), enabling the LLM to autonomously invoke external tools at appropriate points within its chain-of-thought, enabling flexible and knowledge-augmented reasoning.
@@ 2 Related Work
Large language models (LLMs) ensemble methods (Chen et al., 2025a) aim to combine multiple models to leverage their complementary strengths. Existing approaches can be categorized into three paradigms based on integration timing: ensemble-before-inference, ensemble-during-inference, and ensemble-after-inference. Ensemble‐before‐inference methods (Lu et al., 2024a; Ding et al., 2024; Srivatsa et al., 2024; Lu et al., 2024b) first apply a routing mechanism, either pretrained on custom data or trained on the fly, to dispatch each query to the most suitable, specialized model, thereby enabling more cost‐efficient inference. Ensemble-during-inference methods (Huang et al., 2024; Xu et al., 2025b; Park et al., 2025) combine outputs from multiple models at different levels of granularity, including the token level, span level, and reasoning-step level, and then merge the resulting text segments back into the decoding context to iteratively refine the output. Ensemble-after-inference methods (Park et al., 2025; Hu et al., 2025; Du et al., 2024) generate complete responses independently from each candidate LLM and then consolidate them via ranking, majority voting, or fitness scoring to select the highest-quality output for final delivery. In contrast, our CogER learns a lightweight policy network to dispatch each query to a single, optimal inference action, $\mathrm{No\_Think}$, $\mathrm{Think}$, $\mathrm{Extend}$, $\mathrm{Delegate}$, within an MDP, thereby enabling fine-grained per-query adaptation.
Test-Time Scaling (TTS) methods optimize computational resource allocation during inference through adaptive reasoning depth control. Muennighoff et al. (2025) propose budget forcing, a technique to regulate the computation of test time by prematurely stopping the CoT of the model or extending it by repeated insertion of the token ‘wait’ when the model attempts to terminate generation. Aggarwal and Welleck (2025) propose LCPO, a straightforward reinforcement learning (RL) approach designed to maximize accuracy while respecting user‑specified length constraints. Yang et al. (2025b) propose Thinking-Optimal Scaling, which trains the model on seed examples with varied response lengths to learn appropriate reasoning efforts and then self-improves by selecting the shortest correct responses on new tasks. In contrast to TTS methods that only adjust reasoning depth based on coarse difficulty estimates, CogER selects among diverse reasoning modes and seamlessly integrates external tool usage, achieving more versatile resource allocation.
@@ 3 Problem Statement and Motivation
Motivation. In real-world applications, user questions exhibit a wide range of complexity. For example, some queries can be answered in a single step, whereas others require deep, multi-step reasoning or integration of external information sources. However, existing LLMs apply the same reasoning procedure to every query with a high computational cost, which may lead to wasted resources on simple queries and poor performance on complex ones (Sui et al., 2025; Hu et al., 2025).
@@ 4.1 Query Complexity Classification
To efficiently allocate reasoning strategies based on the diverse computational requirements of different queries, we draw inspiration from Bloom’s Taxonomy (Bloom et al., 1956) to classify queries by their cognitive demand. Specifically, we define four levels of query complexity, denoted as $L_{1}$, $L_{2}$, $L_{3}$, and $L_{4}$, with each level representing an increasing degree of reasoning depth and computational demand, as illustrated in Figure 1(a). The details of each level are as follows:
@@ 4.2 Cognitive-Inspired Elastic Reasoning as Markov Decision Process
  • $L_{2}$-level: $\mathrm{Think}$. Apply a chain-of-thought strategy (Wei et al., 2022b) using a moderately sized LLM to produce a concise reasoning trail.
@@ 4.5 Training with Group Relative Policy Optimization
We adopt the Group Relative Policy Optimization (GRPO) (Shao et al., 2024; DeepSeek-AI, 2025) to optimize the parameters $\theta$ of the CogER-Agent due to its superior stability and sample-efficiency.
@@ 5 Experiments
Datasets and Metrics. To train the CogER-Agent, we construct the Reasoning‑Training dataset by randomly sampling 2,000 examples from each of four heterogeneous benchmarks: GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), CommonsenseQA (Talmor et al., 2019), and MedQA (Jin et al., 2020). This unified training set exposes the agent to a wide spectrum of reasoning challenges, from arithmetic word problems to domain‑specific medical questions. For evaluation, we consider both In‑Domain (ID) and Out‑of‑Domain (OOD) settings. ID performance is evaluated on the official test splits of GSM8K, MATH-500, CommonsenseQA, and MedQA, whereas OOD generalization is measured on MAWPS (Koncel-Kedziorski et al., 2016) and CollegeMath (Tang et al., 2024), which are not included in the mixed training set used to fine-tune our CogER-Agent. More details in App. C. We report Exact Match ($EM$) as the metric across all datasets, and record the average parameters (Param.) and Latency used during testing to reflect computational cost.
Baselines. We employ LLMs with varying sizes and architectures, including Qwen2.5-Math-72B-Instruct (Yang et al., 2024b), DeepSeek-R1 (DeepSeek-AI, 2025), DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-14B, and DeepSeek-R1-Distill-Qwen-32B. We also compare against Test-Time Scaling (TTS) methods, including S1 (Muennighoff et al., 2025), L1 (Aggarwal and Welleck, 2025), and ReasonFlux (Yang et al., 2025a).
Implementation Details. In our CogER framework, Qwen2.5-7B-Instruct (Yang et al., 2024a) serves as the CogER-Agent, which assigns queries to appropriate reasoning modules based on their estimated complexity, and also handles all $L_{1}$‑level queries directly. Queries classified as $L_{2}$‑level are escalated to Qwen2.5‑32B‑Instruct for moderate multi‑step reasoning, while $L_{3}$‑level queries are processed by QwQ‑32B (Team, 2025) to support deeper CoT generation. For the most demanding $L_{4}$‑level queries, we invoke our CoTool, whereby QwQ‑32B (Team, 2025) autonomously issues external API calls to enrich its reasoning process. We uniformly capped the generation length at max_token = 8192 for all LLMs. Furthermore, all components are optimized using the AdamW optimizer with a batch size of $24\times 3$ and a learning rate of $5\times 10^{-5}$. The group size $G$ in Eqn. (10) is set to 12. The CogER-Agent is fine-tuned via LoRA with a rank of $r=16$, while all other hyperparameters follow the default settings from the Open-R1 configuration (Face, 2025).
@@ A.1 Large Language Models
Recent advancements in Natural Language Processing (NLP) have culminated in the development of highly capable Large language models (LLMs). These models are predominantly characterized by their utilization of the Transformer architecture (Vaswani et al., 2017) and are pre-trained on vast quantities of textual data. Progressive scaling of model parameters and training datasets has allowed these LLMs to exhibit emergent capabilities (Wei et al., 2022a), demonstrating remarkable proficiency in the handling of a variety of complex tasks. Such tasks include, but are not limited to, high-fidelity question answering (Shao et al., 2023), code generation (Chen et al., 2021), and intermediate-step reasoning (Wei et al., 2022b). Consequently, LLMs have exerted a profound influence on the Artificial Intelligence (AI) community, catalyzing a reevaluation of the prospects for Artificial General Intelligence (AGI) (Zhao et al., 2023). Predicated on their foundational Transformer architecture, extant Large language models can be broadly classified into three primary categories:
@@
Encoder-only models. Encoder-only models, alternatively designated as auto-encoding architectures, are exclusively constructed from encoder modules originating from the Transformer architecture. In such configurations, input data is processed sequentially across multiple layers, facilitating the progressive extraction and encoding of information. A paradigmatic example of this category is BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2019), developed by Google. BERT functions as a language representation model employing bidirectional Transformer encoders. It was pre-trained on a corpus consisting of the BooksCorpus (Zhu et al., 2015) (comprising approximately 800 million words) and English Wikipedia (approximately 2.5 billion words). This pre-training enabled BERT to attain a score of 80.5% on the General Language Understanding Evaluation (GLUE) benchmark and an accuracy of 86.7% on the Multi-Genre Natural Language Inference (MultiNLI) task. Many subsequent encoder-only models are predominantly variants of BERT, including RoBERTa by Meta (Liu et al., 2019) and DeBERTa by Microsoft (He et al., 2021).
@@
Decoder-only models. This category of models is exclusively constructed using decoder modules from the Transformer architecture. Decoder-only models typically implement an auto-regressive mechanism, whereby output sequences are generated on a token-by-token basis. The generation of each token by the decoder is contingent upon the tokens previously generated. Seminal examples in this category include the Generative Pre-trained Transformer (GPT) series (OpenAI, 2023), developed by OpenAI. Illustratively, GPT-3 comprises numerous Transformer decoder layers, featuring up to 175 billion parameters, which established it as one of the most substantial language models at its introduction. This model was pre-trained on a corpus of approximately 300 billion tokens derived from sources such as Common Crawl (Raffel et al., 2020), WebText2, Books1, Books2, and Wikipedia. GPT-3 has demonstrated potent zero-shot and few-shot learning capabilities across a diverse array of language tasks. Beyond the GPT series, a multitude of other decoder-only models have emerged, including OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023a), and Llama 2 (Touvron et al., 2023b) from Meta; PaLM (Chowdhery et al., 2023) and PaLM 2 (Anil et al., 2023) from Google; and BLOOM (Scao et al., 2022) from the BigScience initiative. Furthermore, models such as Qwen 2.5 by Alibaba (Yang et al., 2024a), alongside DeepSeek LLM (Bi et al., 2024) and DeepSeek Coder (Guo et al., 2024) by DeepSeek AI, continue to advance the state-of-the-art in language comprehension, multilingual processing, and domain-specific generation.
@@
Encoder-decoder models. This architectural class integrates both encoder and decoder modules from the Transformer framework. Such models aim to amalgamate the respective strengths of the aforementioned architectures, thereby effectively addressing tasks that necessitate comprehensive input understanding coupled with the generation of extended output sequences. Prominent extant encoder-decoder models encompass GLM from Tsinghua University (Du et al., 2022); T5 (Raffel et al., 2020), FLAN-T5 (Chung et al., 2024), and UL2 (Tay et al., 2023) from Google; and BART (Lewis et al., 2020) from Meta. For instance, the GLM model employs an autoregressive blank infilling objective. This methodology is designed to tackle three fundamental challenges in NLP: natural language understanding (NLU), unconditional text generation, and conditional text generation. With a reported maximum of 130 billion parameters, GLM was pre-trained on a corpus including BookCorpus (Tay et al., 2023) and Wikipedia. GLM has demonstrated superior performance over BERT on the SuperGLUE benchmark, exhibiting an improvement of 4.6%–5.0%. Moreover, it significantly surpasses FLAN-T5 in both NLU and generation tasks, even when utilizing fewer parameters and less training data.